Linear Algebra (Introduction)​6. Prove the following identity
Answer #1

α is a scalar quantity. Then

α (A+B)= α A +α B by the Linearity property.

Let α=2, A=3, B=5

α (A+B)=2(3+5)=16

α A+α B= (2*3)+(2*5)=16

Hence Linearity property satisfied.